OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x + 3y = –12 A. x-intercept –6, y-intercept –4 B. x-intercept 4, y-intercept –6 C. x-intercept –6, y-intercept 4 D. x-intercept –4, y-intercept 6

OpenStudy (anonymous):

plug in 0 for each y and x, one at a time. 2(0) + 3y = –12 2x + 3(0) = –12 then divide.

OpenStudy (anonymous):

Ok, let me try that! @silverr

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

The x-intercept is -6 and the y-intercept is -4. Is that correct? :)

OpenStudy (anonymous):

@silverr

OpenStudy (anonymous):

lets see 2(0) + 3y = –12 -12 / 3 = -4 y = -4 2x + 3(0) = –12 -12 / 2 = -6 x = -6 you are correct !

OpenStudy (anonymous):

Yay! XD Thanks for you help!

OpenStudy (anonymous):

no problem :) i learned this recently, so it's still fresh.

OpenStudy (anonymous):

Hehe :)